OpenStudy (magbak):

I WILL AWARD MEDAL !!!!!!!!!!!!!!!!!!!! PLEASE HELP!!!!!!!!!!!!!!!!! Using complete sentences, describe the benefit of simplifying rational expressions before multiplying.

OpenStudy (magbak):

@tcarroll010 @jim_thompson5910 @whpalmer4 @Hero

hero (hero):

Well, it makes computations as simple as possible that would otherwise be more expanded than necessary.

hero (hero):

Simplifying before computing leads to more simpler expressions. Not simplifying before computing leads to more complex expressions. Therefore, your first impulse when solving any kind of equation or reducing any expression should be to 'simplify' first if possible.

OpenStudy (magbak):

Also can you help me in this question also Using complete sentences, describe how you would check the quotient of two rational expressions. Create and demonstrate your own example of this concept.

hero (hero):

Are you taking a test?

hero (hero):

Be honest.

hero (hero):

A quiz is essentially a test, just on a smaller scale.

OpenStudy (magbak):

Yes

hero (hero):

@tkhunny, I figured you were shadowing me.

OpenStudy (tkhunny):

"Using complete sentences, describe the benefit of simplifying rational expressions before multiplying. " Here's mine: There may or may not be benefits to simplifying rational expressions before multiplying. If you don't happen to need the simplified version, then it was a waste of time.

hero (hero):

Example @tkhunny

hero (hero):

I've always simplifed first before simplifying expressions. I have yet to 'not need' the simpler version.

hero (hero):

Although there was that one problem I did....

OpenStudy (magbak):

Help me with this one please Using complete sentences, describe how you would check the quotient of two rational expressions. Create and demonstrate your own example of this concept.

OpenStudy (tkhunny):

Fair enough: \(\dfrac{y^{2} - 2y}{y} = 0\) -- Solve for \(y^{2}\)

OpenStudy (magbak):

The entire question please >1/2

hero (hero):

If you're solving for \(y^2\) then you're solving for \(y^2\). It would make no sense to simplify in this case.

hero (hero):

That's why we say simplify (if necessary) usually.

OpenStudy (tkhunny):

So, you're buying it? :-)

hero (hero):

In this case, it isn't necessary

OpenStudy (magbak):

Oh so you are trying to trick me huh.

OpenStudy (magbak):

@tkhunny I sm talking to you not @Hero

hero (hero):

However, suppose someone did decide to simplify first. They would get y - 2 = 0 y = 2 Then multiply both sides by y to get y^2 = 2y So...same difference

OpenStudy (tkhunny):

@magbak Just trying to be thorough. You'll be the only one in your class with a counter example to the original premise. :-) @hero For sure. Unique answers don't care how you find them?

OpenStudy (magbak):

I was hoping to get a answer to my question not to get lost in a mathematical argument

hero (hero):

This isn't a mathematical argument per se.

OpenStudy (magbak):

It is not so where is my answer in all of this. I am not trying to be rude or anything.

OpenStudy (tkhunny):

Okay, without the finer points: Simplification prior to multiplication keeps all expressions smaller and more manageable. How's that?

OpenStudy (magbak):

Is this for the second question or the main post I already understood the main question not the second question.

OpenStudy (tkhunny):

On the other hand: \(\dfrac{x^{2}-1}{x+1}\cdot\dfrac{x-1}{x^{2}-1}\) I would NOT simplify that one factor at a time.

hero (hero):

What do you mean by that @tkhunny?

OpenStudy (tkhunny):

\(\dfrac{x^{2} - 1}{x^{2} - 1}\cdot\dfrac{x-1}{x+1} = \dfrac{x-1}{x+1}\) Much better way to go.

hero (hero):

That's still simplification in my opinion. You didn't expand or multiply.

OpenStudy (tkhunny):

The premise is "simplifying rational expressions before multiplying. " I eliminated the multiplication (or reduced the multiplicand to unity.)

hero (hero):

Yet another reason for simplifying. In cases such as this, you may not even have to multiply anything.

OpenStudy (tkhunny):

@magbak " Create and demonstrate YOUR OWN example of this concept. " (bold added) No one can help you with this.

OpenStudy (magbak):

WHY NOT YOU ARE SO LAYZY

hero (hero):

And you're not @magbak?

OpenStudy (tkhunny):

I'm not in your head.

OpenStudy (magbak):

I am just joking I am probably the most laziest person on earth after a couch potato.

OpenStudy (tkhunny):

Joke accepted. Now, get to work on that example.